# Using Large Language Models for Expert Prior Elicitation in Predictive Modelling

### v2

@@ A.9 Dataset memorisation
To better contextualise the use of in-context learning in the tested datasets, we will use the method proposed in Bordt et al. [42] to quantify the extent to which the language model already knows information about the dataset.
Table 3 shows the results of the header completion test described in Bordt et al. [42], which tests whether the language model has memorised the first few rows of the dataset. The column names along with a given number of the first rows are given to the language model as a string, with the final of these rows truncated. The language model is then asked to generate the next $500$ tokens.
We use the Levenshtein distance [51, 52], a string matching metric, to calculate the agreement between the true header completion and the one provided by the language model. The Levenshtein distance counts the number of single-character changes required to transform one string into another, only allowing for insertions, deletions, and substitutions. The value can take a minimum value of 00 – suggesting that the strings are equal and a maximum value of the greatest length of the two strings.
@@
Where $\text{Levenshtein}(\cdot,\cdot)$ calculates the Levenshtein distance between two strings, $\text{min}\{\cdot,\cdot\}$ returns the minimum value of the inputs, and $|s|$ refers to the length of $s$. This new value ranges between 00 and $1$, with 00 indicating a perfect match between the strings and $1$ meaning that they were different at all positions.
@@
Breast Cancer California Housing Wine Quality Heart Disease
0.00 0.15 0.21 0.00
The values in Table 3 demonstrate that the language model has completely memorised the first few rows of the Breast Cancer and Heart Disease datasets, demonstrated by a normalised Levenshtein distance value of 0. This could be because the first few rows of these datasets commonly appear within the language model’s training data. For California Housing and Wine Quality, we see some level of memorisation, with the language model being able to reproduce a large amount of the true header, however, to a much lesser extent than what we see for Breast Cancer and Heart Disease. As we might expect, when we perform the header test on the synthetic dataset generated using the relationship $y=2x_{1}-x_{2}+x_{3}$, we get a normalised Levenshtein distance of $0.7$, which shows a large difference between the true header string and the completion from the language model. The UTI task has not been tested, as it contains sensitive information and is currently a private dataset.
@@
Breast Cancer California Housing Wine Quality Heart Disease
0.43 ± 0.13 0.58 ± 0.04 0.31 ± 0.04 0.33 ± 0.05
We supplement these results with those of Table 4, which shows the results of the row completion test. Here we calculate the normalised Levenshtein distance of the true completions and the language model’s completion of a random row from the dataset. In these tests, the language model is provided with $10$ rows (in order from the original dataset) and is asked to complete the following row. The values in Table 4 represent the mean and standard deviation of the test using $25$ random rows. Here, in contrast to the header test, we see that the language model has not memorised all rows from the dataset. This is because the language model’s completion differs from the true row completion in all cases to varying degrees.
This is likely because when the datasets appear within the language model’s training data, their column names and initial rows are given, but not the entire dataset.
@@ A.9.1 How does this affect our results?
Further, since for prior elicitation we only prompt the language model for a prior distribution of a predictive model given the feature names and a description of the dataset, the memorisation of rows from the data is unlikely to have had an impact on the results. In addition, we see that elicited priors provide improved predictions over the uninformative prior for the private or synthetic datasets, as well as the public datasets.
@@ A.10 Task descriptions with more information
To understand how the task description affects the performance of the priors elicited from the language model, we studied the synthetic task, $y=2x_{1}-x_{2}+x_{3}$, and described the relationship between features and targets in varied detail. We additionally test an uninformative prior to contextualise the results. For each of the levels of detail shown in Figure 13, we included the following text in the task description that was provided to the language model:
@@
For each prior elicited using these levels of detail, in Figure 13, the corresponding $\log_{10}\text{MSE}$ on the test set can be seen for a linear regressor after observing the given number of data points. This figure illustrates the effects of varied descriptions and shows that the improvements made to the MSE diminish as the task description contains more detail. The biggest improvement came from providing the language model with a single feature relationship – enabling an MSE that is two orders of magnitude better.
@@
However, what is unexpected is that for this task, using an uninformative prior ($\theta\sim\mathcal{N}(\theta|0,1)$) provides better posterior performance than specifying to the language model that the relationship between the target and features is linear (Figure 13). This is explained by inspecting the prior distribution elicited when given this information, which shows that the language model responded with a similar mean and standard deviation for a Gaussian for each of the $100$ variations of the prompts:
@@
An expert interacting with the language model might be able to provide further information about a task that improves the elicited priors. In Figure 14 we show the parameter distributions elicited from the language model for the different descriptions we provide. This figure shows that, as we provide additional information about each feature, the language model responds by providing a prior distribution that reflects its updated knowledge of the task. This demonstrates the potential of using an expert to construct more detailed task descriptions, which the language model can translate into parameter distributions and supplements the results in Section 3.2.
@@ A.11 How many ways should we describe a task?
In this section we look at how many of the task descriptions are required to be able to get a similar prior distribution to that elicited within our experiments, using $100$ task descriptions.
@@
It is interesting to see that Breast Cancer benefitted most from increasing the number of task descriptions, since this dataset contains the largest number of features. It suggests that more task descriptions are required for those datasets with high-dimensional data. This makes intuitive sense, since we would expect to require a larger mixture of Gaussian components to better describe the prior over for a higher-dimensional dataset.
This figure shows that as the number of task descriptions increases, we see smaller changes in the energy distance, and reduced variability in the samples. This suggests that increasing the number of task descriptions provides diminishing returns.
@@ A.13.1 Other OpenAI models
Figure 16: Test accuracy of the posterior distribution for different OpenAI language models. Here we show the results when using priors elicited from two other OpenAI language models and an uninformative prior $\mathcal{N}(\theta|0,1)$ on a linear model. These results are calculated on a test set of the datasets after each model is trained on the given number of data points. These values are the result of $5000$ sampled parameters and $10$ random splits of the training and testing data. For California Housing, a regression task, we report the mean squared error on log-scale. This figure shows the line plot of the average mean posterior accuracy or mean squared error (MSE), with the error bars representing the 95% confidence interval. The green arrows point in the direction of improvement in the metric.
In Figure 16 we also present the posterior results of the priors elicited from both GPT-4-turbo and GPT-4o-mini. GPT-4-turbo incurs a significantly higher cost than the GPT-3.5-turbo model used for the main results in this paper, whilst GPT-4o-mini provides a less expensive option.
Specifically, to elicit all priors presented in Figure 16, GPT-4o-mini cost \$$\$0.32$ (on average \$$\$0.053$ per dataset), whilst GPT-4-turbo cost \$$\$7.56$ (on average \$$\$1.26$ per dataset). This is compared to the cost of \$$\$0.41$ (on average \$$\$0.068$ per dataset) for GPT-3.5-turbo (presented in A.12). The more features a dataset contains, the more expensive priors are to elicit since the number of input and output tokens is greater.
Figure 16 demonstrates that in most cases, GPT-4-turbo provides priors that are significantly better than GPT-3.5-turbo but falls behind in the regression tasks. On Breast Cancer, UTI, and Wine Quality, the priors elicited from GPT-4-turbo allow for a significantly better posterior accuracy, and unlike GPT-3.5-turbo, provide an informative prior for Heart Disease. However, for both regression tasks (the synthetic task and the California Housing task), the GPT-3.5-turbo priors allow for a better mean squared error (MSE) than GPT-4-turbo.
GPT-4o-mini often produces prior distributions that are equal to or worse than GPT-3.5-turbo. For the synthetic task, the priors elicited from GPT-4o-mini are as good as an uninformative prior; for Breast Cancer and UTI, they are as good as GPT-3.5-turbo; for Wine Quality, Heart Disease, and California Housing however, they are worse than GPT-3.5-turbo. For GPT-4o-mini, eliciting priors for Heart Disease caused significant problems, as they were worse than both the uninformative prior and those elicited from GPT-3.5-turbo.
@@ A.13.2 Open source language models
Figure 17: Test accuracy of the posterior distribution for Llama language models. Here, we show the results when using priors elicited from two of the Llama language models and an uninformative prior $\mathcal{N}(\theta|0,1)$ on a linear model. These results are calculated on a test set of the datasets after each model is trained on the given number of data points. These values are the result of $5000$ sampled parameters and $10$ random splits of the training and testing data. For California Housing, a regression task, we report the mean squared error on log-scale. This figure shows the line plot of the average mean posterior accuracy or mean squared error (MSE), with the error bars representing the 95% confidence interval. The green arrows point in the direction of improvement in the metric.
@@
In contrast to the OpenAI models in Appendix A.13.1, these models are publicly available and open source after being granted access. This means that they can be used without the need for connecting to an online API, and when locally hosted, all data and inference can be performed on device, enabling the use of our framework within a closed and secure environment.
@@
The experiments presented in this section use the same task descriptions as given to GPT-3.5-turbo, except that because Llama models often fail to return a JSON structure, we suffix the prompt with an additional instruction to return a JSON. When sampling the posterior with the elicited priors, we use the same methods and options as given in A.5.2.
@@
Figure 17 illustrates the inconsistent posterior accuracy of the priors elicited from the Llama family of tested language models. These models are significantly smaller in capacity to GPT-3.5-turbo, justifying their inconsistent elicited priors, but in some of the datasets, they still outperform the uninformative prior.
@@
Specifically for Breast Cancer, the 3B parameter Llama model provides an intermediate accuracy between an uninformative prior and those elicited from GPT-3.5-turbo, and for the synthetic task $y=2x_{1}-x_{2}+x_{3}$ and California Housing, the 8B parameter Llama provides priors of equal quality to GPT-3.5-turbo. However, for the UTI, Wine Quality, and Heart Disease datasets, both Llama models provide prior distributions that are worse quality than the uninformative prior. It is possibly not surprising that the Llama-3.2-3B-Instruct provided priors that were equal or worse than Llama-3.1-8B-Instruct in all cases except for predicting Breast Cancer since these models are significantly smaller in size.
@@
Llama-3.1-70B-Instruct-BF16 is more competitive, sometimes providing priors that allow for better posterior distributions than GPT-3.5-turbo, but on other occasions, providing priors with similar posterior performance to the uninformative prior. On the UTI and Heart Disease datasets, the prior distributions elicited from Llama-3.1-70B-Instruct-BF16 provide better posterior accuracy than those from GPT-3.5-turbo. Surprisingly, on Heart Disease prediction in particular, priors from the 70B Llama model provided a posterior accuracy greater than GPT-4-turbo. Additionally, when predicting Breast Cancer, priors form Llama-3.1-70B-Instruct-BF16 achieved the same posterior accuracy as GPT-3.5-turbo for all of the number of training examples tested. These three datasets suggest that Llama-3.1-70B-Instruct-BF16 can be a useful model for eliciting prior distributions for a predictive task. However, when eliciting priors for the Wine Quality, California Housing, and $y=2x_{1}-x_{2}+x_{3}$ tasks, we found that Llama-3.1-70B-Instruct-BF16 provided distributions that were not competitive with GPT-3.5-turbo.
@@
Qwen-2.5-14B-Instruct is similarly inconsistent in the posterior performance that the priors provide. For Heart Disease, it achieves a greater posterior accuracy than all models tested, except for the 70B Llama model, Llama-3.1-70B-Instruct-BF16. For the UTI task, after seeing $10$ demonstrations, it performs comparably to Llama-3.1-70B-Instruct-BF16, and better than GPT-3.5-turbo. On Wine Quality it achieves the greatest posterior accuracy after $20$ observed data points. However, on the California Housing, Breast Cancer, and the synthetic task, it performs worse than GPT-3.5-turbo, but is likely smaller in capacity. Overall, Qwen-2.5-14B-Instruct does not provide consistent performance across tasks, but has the benefit of performing similarly to Llama-3.1-70B-Instruct-BF16 on many tasks with a significantly smaller compute requirement.

### v3

@@ A.9 Dataset memorisation
To better contextualise the use of in-context learning in the tested datasets, we will use the method proposed in Bordt et al. [42] to quantify the extent to which the language model already knows information about the datasets.
We start with the header completion test, which checks whether the language model has memorised the first few rows of the dataset. Here, the column names along with a given number of the first rows are given to the language model as a string, with the final of these rows truncated. The language model is then asked to generate the next $500$ tokens.
We then use the Levenshtein distance [51, 52], a string matching metric, to calculate the agreement between the true header completion and the one provided by the language model. The Levenshtein distance counts the number of single-character changes required to transform one string into another, only allowing for insertions, deletions, and substitutions. The value can take a minimum value of 00, suggesting that the strings are equal, and a maximum value of the longest length of the two strings.
@@
Here, $\text{Levenshtein}(s_{1},s_{2})$ calculates the Levenshtein distance between two strings, and $\text{min}\{|s_{1}|,|s_{2}|\}$ returns the minimum length of the strings. This new value ranges between 00 and $1$, with 00 indicating a perfect match between the strings and $1$ implying that they were different at all positions.
@@
Heart Disease Diabetes Hypothyroid Breast Cancer
0.00 0.21 0.22 0.00
The values in Table 3 demonstrate that the language model has completely memorised the first few rows of the Breast Cancer and Heart Disease datasets, demonstrated by a normalised Levenshtein distance value of 0. This could be because the first few rows of these datasets commonly appear within the language model’s training data. For the Diabetes and Hypothyroid datasets, we see some level of memorisation, with the language model being able to reproduce a large amount of the true header. As we might expect, when we perform the header test on the synthetic dataset generated using the relationship $y=2x_{1}-x_{2}+x_{3}$, we get a normalised Levenshtein distance of $0.7$, which shows a large difference between the true header string and the completion from the language model. The UTI task has not been tested, as it is a private dataset.
@@
Heart Disease Diabetes Hypothyroid Breast Cancer
0.33 ± 0.05 0.38 ± 0.05 0.20 ± 0.05 0.43 ± 0.13
We supplement these results with those of Table 4, which shows the results of the row completion test. Here we calculate the normalised Levenshtein distance of the true completions and the language model’s completion of a random row from the dataset. In these tests, the language model is provided with $10$ rows (in order from the original dataset) and is asked to complete the following row. The values in Table 4 represent the mean and standard deviation of the test using $25$ random rows. Here, in contrast to the header test, we see that the language model has not memorised all rows from the dataset. This is because the language model’s completion differs from the true row completion in all cases.
We likely see these results because when the datasets appear within the language model’s training data, their column names and initial rows are usually given, but not the entire dataset.
@@ A.9.1 How does this affect our results?
Further, since for prior elicitation we only prompt the language model for a prior distribution of a predictive model given the feature names and a description of the dataset, the memorisation of rows from the data is unlikely to have had an impact on these results. In addition, we see that elicited priors provide improved predictions over the uninformative prior for the private and synthetic datasets, as well as the public datasets.
@@ A.10 Task descriptions with more information
To understand how the detail of the task description affects the performance of the priors elicited from the language model, we studied the synthetic task, $y=2x_{1}-x_{2}+x_{3}$, and described the relationship between features and targets in varied detail. We additionally test an uninformative prior to contextualise the results. For each of the levels of detail shown in Figure 14, we included the following text in the task description that was provided to the language model:
@@
For each prior elicited using these levels of detail, in Figure 14, the corresponding $\log_{10}\text{MSE}$ on the test set can be seen for a linear regressor after observing the given number of data points. This figure illustrates the effects of varied descriptions and shows that the improvements made to the MSE diminish as the task description contains more detail. The biggest improvement came from providing the language model with a single feature relationship – enabling an MSE that is two orders of magnitude better. This suggests that it is most helpful to provide more detail when the language model initially knows very little about a task.
@@
However, what is unexpected is that for this task, using an uninformative prior ($\theta\sim\mathcal{N}(\theta|0,1)$) provides better posterior performance than specifying to the language model that the relationship between the target and features is linear (Figure 14). This is explained by inspecting the prior distribution elicited when given this minimal information, which shows that the language model responded with a similar mean and standard deviation for a Gaussian for each of the $100$ variations of the prompts:
@@
An expert interacting with the language model might be able to provide further information about a task that improves the elicited priors. In Figure 15 we show the parameter distributions elicited from the language model for the different descriptions we provide. This figure shows that, as we provide additional information about each feature, the language model responds by providing a prior distribution that reflects its updated knowledge of the task. This demonstrates the potential of using an expert to construct more detailed task descriptions, which the language model can translate into parameter distributions. This experiment provides a controlled setting that supplements the results in Section 3.2.
@@ A.11 How many ways should we describe a task?
In this section, we look at how many of the task descriptions are required to be able to get a similar prior distribution to that elicited within our experiments, using $100$ task descriptions.
@@
It is interesting to see that Breast Cancer benefitted most from increasing the number of task descriptions, since this dataset contains the largest number of features. It suggests that more task descriptions are required for those datasets with high-dimensional data. This makes intuitive sense since we would expect to require a larger mixture of Gaussian components to better describe the prior over for a higher-dimensional dataset.
This figure shows that as the number of task descriptions increases, we see smaller changes in the energy distance and reduced variability in the samples. This suggests that increasing the number of task descriptions provides diminishing returns.
@@ A.13.1 Other OpenAI models
Figure 17: Test accuracy of the posterior distribution for different OpenAI language models. Here we show the posterior results when using priors elicited from two other OpenAI language models and an uninformative prior $\mathcal{N}(\theta|0,1)$. These results are calculated on a test set of the datasets after each model is trained on the given number of data points. The values are the result of $5000$ sampled parameters for each of $5$ chains, and $10$ random splits of the training and testing data. For the regression tasks, we report the mean squared error on log-scale. This figure shows the line plot of the average mean posterior accuracy or mean squared error over the $10$ splits, with the error bars representing the 95% confidence interval. The green arrows point in the direction of improvement in the metric.
In this section we compare the posterior results of priors elicited from GPT-3.5-turbo to GPT-4-turbo which incurs a significantly higher cost, and GPT-4o-mini which provides a less expensive option.
Specifically, to elicit all priors presented here, GPT-4o-mini cost \$$\$0.32$ (on average \$$\$0.053$ per dataset), whilst GPT-4-turbo cost \$$\$7.56$ (on average \$$\$1.26$ per dataset). This is compared to the cost of \$$\$0.41$ (on average \$$\$0.068$ per dataset) for GPT-3.5-turbo (presented in A.12). As a rule, the more features a dataset contains or the longer the task description, the more expensive priors are to elicit since the number of input and output tokens is greater.
Figure 17 shows that in most cases, GPT-4-turbo provides priors that are significantly better than GPT-3.5-turbo but falls behind in the synthetic and Hypothyroid prediction tasks. On the Breast Cancer, UTI, and Heart Disease tasks, the priors elicited from GPT-4-turbo allow for a greatly improved posterior accuracy, and unlike GPT-3.5-turbo, provide an informative prior for Heart Disease. In addition, on the Diabetes predictive task, GPT-4-turbo achieves similar mean squared error to GPT-3.5-turbo, with a small improvement in performance when only a few training data points are available.
On the other hand, GPT-4o-mini often produces prior distributions that are equal to or worse than GPT-3.5-turbo in terms of their posterior performance in all tasks except one. For the synthetic, Diabetes, and Heart Disease prediction tasks, GPT-4o-mini provides prior distributions that achieve similar or worse performance to the uninformative prior. However, for the UTI and Breast Cancer predictive tasks they achieve similar posterior accuracy to GPT-3.5-turbo at a lower price point. Also, on the Hypothyroid dataset GPT-4o-mini provides prior distributions that achieve greater accuracy than those from GPT-3.5-turbo and GPT-4-turbo for all training sizes, whilst requiring significantly less cost.
@@ A.13.2 Open source language models
Figure 18: Test accuracy of the posterior distribution for Llama language models. Here, we show the results when using priors elicited from two of the Llama language models and an uninformative prior $\mathcal{N}(\theta|0,1)$ on a linear model. These results are calculated on a test set of the datasets after each model is trained on the given number of data points. The values are the result of $5000$ sampled parameters for each of $5$ chains, and $10$ random splits of the training and testing data. For the regression tasks, we report the mean squared error on log-scale. This figure shows the line plot of the average mean posterior accuracy or mean squared error over the $10$ splits, with the error bars representing the 95% confidence interval. The green arrows point in the direction of improvement in the metric.
@@
In contrast to the OpenAI models in Appendix A.13.1, these models are publicly available and open source. This means that they can be used without the need for connecting to an online API, and when locally hosted, all data and inference can be performed on device, enabling the use of our framework within a closed and secure environment.
@@
The experiments presented in this section use the same task descriptions as given to GPT-3.5-turbo, except that because these smaller models sometimes fail to return a JSON structure, we suffix the prompt with an additional instruction to return a JSON. When sampling the posterior with the elicited priors, we use the same methods and options as given in A.5.2.
@@
Figure 18 illustrates the inconsistent posterior accuracy of the priors elicited from the Llama family of tested language models. These models are significantly smaller in capacity to GPT-3.5-turbo, justifying their inconsistent elicited priors, but in some of the datasets, they still produce the greatest posterior performance.
@@
Specifically for Breast Cancer, the 3B parameter Llama model provides an intermediate accuracy between an uninformative prior and those elicited from GPT-3.5-turbo, and for the synthetic and Hypothyroid tasks, the 8B parameter Llama provides priors of equal quality to GPT-3.5-turbo. However, for the UTI and Heart Disease datasets, both the 3B and 8B Llama models provide prior distributions that are of worse quality than even the uninformative prior. It is possibly not surprising that the Llama-3.2-3B-Instruct provided priors that were equal or worse than Llama-3.1-8B-Instruct in all cases except for predicting Breast Cancer since these models are significantly smaller in size.
@@
Llama-3.1-70B-Instruct-BF16 is more competitive, sometimes providing priors that allow for better posterior distributions than GPT-3.5-turbo, but on other occasions, providing priors with similar posterior performance to the uninformative prior. On the UTI and Heart Disease datasets, the prior distributions elicited from Llama-3.1-70B-Instruct-BF16 provide better posterior accuracy than those from GPT-3.5-turbo. Surprisingly, on Heart Disease prediction in particular, priors from the 70B Llama model provided a posterior accuracy greater than GPT-4-turbo. Additionally, when predicting Breast Cancer, priors form Llama-3.1-70B-Instruct-BF16 achieved the same posterior accuracy as GPT-3.5-turbo for all of the number of training examples tested. These three datasets suggest that Llama-3.1-70B-Instruct-BF16 can be a useful model for eliciting prior distributions for a predictive task. However, when eliciting priors for the Diabetes and synthetic tasks, we found that Llama-3.1-70B-Instruct-BF16 provided distributions that were not competitive with GPT-3.5-turbo.
@@
Qwen-2.5-14B-Instruct is similarly inconsistent in the posterior performance that the priors provide. For Heart Disease and Hypothyroid prediction, it achieves a greater posterior accuracy than all models in Figure 18, except for the 70B Llama model, Llama-3.1-70B-Instruct-BF16. For the UTI task, after seeing $10$ demonstrations, it performs comparably to Llama-3.1-70B-Instruct-BF16, and better than GPT-3.5-turbo. However, on the Breast Cancer and synthetic tasks, it performs worse than GPT-3.5-turbo, but is likely smaller in capacity. Overall, Qwen-2.5-14B-Instruct provides inconsistent performance across tasks, but has the benefit of performing similarly to Llama-3.1-70B-Instruct-BF16 on many tasks with a significantly smaller compute requirement.